OpenStudy (anonymous):

hello i need serious help here 5+4/9/5-4/5

OpenStudy (anonymous):

193/45...i think

OpenStudy (anonymous):

amk11 that isnt right it says find the quotient

OpenStudy (anonymous):

ill become your fan because you tried

OpenStudy (amistre64):

its important to know which order to do this in. For example: 4/2/2 has 2 different answers. (4/2)/2 = 2/2 = 1 4/(2/2) = 4/1 = 4

OpenStudy (anonymous):

???????????? the problem is 5+ 4/9 over 5-4/5 and then it says simplify.

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

wrong thing

OpenStudy (amistre64):

Ill assume then: 5 + (4/9)/5 - (4/5) ?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so far I get 5 - (32/45)....thinking

OpenStudy (anonymous):

5 + ((4 / 9) over 5) - (4 / 5) = 4.28888889

OpenStudy (amistre64):

193/45 is what I get.... might need to reduce

OpenStudy (anonymous):

35/27

OpenStudy (anonymous):

was the answer...

OpenStudy (anonymous):

i got it wrong....

OpenStudy (anonymous):

5+(2/5)/3-(5/7)

OpenStudy (anonymous):

5+4/35-36/45

OpenStudy (anonymous):

do the fractions

OpenStudy (anonymous):

5-31/45

OpenStudy (anonymous):

mathtio thats your job...

OpenStudy (anonymous):

mathtio you got it wrong...thanks alot

OpenStudy (amistre64):

(77/15) - (5/7)....

OpenStudy (anonymous):

225/45-31-45

OpenStudy (anonymous):

????? im lost

OpenStudy (anonymous):

sorry...225/45-31/45

OpenStudy (amistre64):

464/105....give or take :)

OpenStudy (anonymous):

194/45

OpenStudy (anonymous):

reduce

OpenStudy (amistre64):

reduce, reuse, replace :)

OpenStudy (anonymous):

i got it wrong...again

OpenStudy (anonymous):

????

OpenStudy (anonymous):

????

OpenStudy (anonymous):

arghhh confused, rewrite the question

OpenStudy (anonymous):

you should check out this website:WOLFRAMALPHA.COM

OpenStudy (anonymous):

just punch in equations and it will solve them.

OpenStudy (anonymous):

best thing since slice bread.

OpenStudy (anonymous):

Just solve the division first then operate the subtraction.